Circular No. 6636

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP) Z. Sekanina, Jet Propulsion Laboratory, reports: “Available informa-

tion on the sodium tail (IAUC 6631, 6634), combined with the photoionization lifetime of 0.56 day for sodium atoms at r = 0.97 AU from the sun (inferred as an upper limit to their destruction lifetime; Spinrad and Miner 1968, Ap.J. 153, 355) and with the radiation-pressure acceleration of 52 cm/s^{2} at the same r, or β = 83 (Wurm 1963, Icarus 2, 29), shows that sodium atoms could not be released directly from the nucleus or from dust grains of any size (visible or invisible). The first option is eliminated because of the short destruction lifetime, the second — in addition — because the accelerations on all particulates in comets are β_{dust} < 3, more than one order of magnitude lower than required. Thus, a sodium tail of this kind would in the best case trace the leading boundary of the dust tail. The most probable parents consistent with observation are sodium-bearing molecular species with relatively long dissociation lifetimes, which could have been initially released from the comet in dust particles but separated soon after their ejection. The observed width of the sodium tail is consistent with the above photoionization lifetime at an expansion velocity of < 7 km/s. If this velocity is provided entirely by the parent’s dissociation, the implied dissociation energy is several eV, and Spinrad and Miner list a number of molecules that would generally fit this constraint. The combined imaging and spectroscopic observations from Apr. 19.90 require β_{parent} = 91 (coincident within errors with the sodium’s β value) and the parent’s dissociation lifetime of τ_{parent} = 2.2 days to explain the tail’s presence in p.a. 56^{o} at a distance of 1^{o}.5 from the nucleus. The calculated line-of-sight velocity is +114 km/s, close to the observed value. On Apr. 20.85 UT at 0^{o}.7 from the nucleus, the implied τ_{parent} = 1.5 days (for the same value of β_{parent}) and the line-of-sight velocity is calculated to be 57 km/s after correction for the comet’s contribution, in excellent agreement with observation. For Apr. 16.88, the tail’s length of 6^{o}.6 would require τ_{parent} = 4.8 days for β = 91, but the calculated p.a. is then 47^{o}. A better fit is provided by β ≃ 450 at τ_{parent} = 2.2 days, in which case p.a. = 51^{o}, in good agreement with observation. In any case, the constraint implied by the tail’s maximum length of 6^{o}.6 is (τ\sqrt{β})_{parent} ≤ 50r^{2} days, which could be tested by a continuing monitoring of the tail’s appearance (after accounting for production variations with r).”

1997 April 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green